OpenStudy (anonymous):

7.7^4x+3=10.2 solve for x?

OpenStudy (anonymous):

Assuming you mean \[ 7.7^{4x+3} = 10.2 \] the answer is approximately -0.4655638116.

OpenStudy (anonymous):

Assuming you mean \[ 7.7^{4x}+3=10.2 \] the answer is approximately 0.2417770632.

OpenStudy (anonymous):

Assuming you mean \[ 7.7^4\cdot x+3=10.2 \] the answer is approximately 0.002048186955.

OpenStudy (anonymous):

\[7.7^{4x}+3=10.2\]\[7.7^{4x}=7.2\]\[\ln(7.7^{4x})=\ln(7.2)\]\[4xln(7.7)=\ln(7.2)\]\[x=\frac{\ln(7.2)}{4\ln(7.7)} = 0.241777\]

OpenStudy (anonymous):

The first two you can find by taking the logarithm.

OpenStudy (anonymous):

Thank you sooo much.